OpenStudy (anonymous):

find the exact value of sec theta without finding theta if cot theta =-5 and sin theta>0

OpenStudy (anonymous):

we have \[\cot\theta={1\over\tan\theta}\\ \sec^2\theta=1+\tan^2\theta \]

OpenStudy (anonymous):

you'd get two answers for sec theta one positive and one negative. from that fact that "cot is negative" and "sin is positive", what should be "sec?"

OpenStudy (anonymous):

positive??

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

im not really sure how to start the problem